A garden supply store sells two types of lawn mowers. Total sales of mowers for the year were $8379.70. The total number of mowers sold was 30. The small mower costs $249.99. The large mower costs $329.99. Find the number sold of each type of mower.
Found 2 solutions by richwmiller, gsmani_iyer:
Answer by richwmiller(9143)

s+l=30
249.99s+329.99l=8379.70
11 large and 19 small

Answer by gsmani_iyer(180)


The total No. of mowers = 30

Let the No. of small mowers be = x; So,
the No. of large mowers = (30-x)
Sale of small mowers = ($249.99)(x) and the
Sale of large mowers = $329.99(30-x)
= $9899.70 - 329.99x
Total sale = $249.99x + $9899.70 - 329.99x = $8379.70
= -80x + $9899.70 = $8379.70. (By transposing,
$9899.70 - $8379.70 = 80x
= $1520.00 = 80x. Therefore,
= x = $1520/80 = 19 small mowers.
So the large mowers (30 - x) = 30 -19 =11 large mowers.
Small mowers = 19; Large mowers = 11. Answer.
I think you have understood this. If not, feel free to contact me anytime for any doubts in maths online.
God Bless you.
gsm